Right an equation of a line through the point (5,2) and parallel to 3x +4y=-50. Write your answer in. Point slope form
Answer by Alan3354(69443)

Right [sic] an equation of a line through the point (5,2) and parallel to 3x +4y=-50. Write your answer in. Point slope form
----------
Step 1, find the slope,m, of the line.
Step 2, use y-y1 = m*(x-x1) where (x1,y1) is the point.